My daughter has this critical thinking problem and i'm a bit rusy with my algebra :( Can you help?
"An apple costs the same as 2 oranges. Together, an orange and a banana cost .10 cents more than an apple. Two oranges cost .15 cents more than a banana. What is the cost for one of each fruit? "

Answer by stanbon(75887)

"An apple costs the same as 2 oranges. Together, an orange and a banana cost .10 cents more than an apple. Two oranges cost .15 cents more than a banana. What is the cost for one of each fruit? "
----------------------------
Equations:
a = 2r
r + b = a + 10
2r = b + 15
--------------------
Rearrange the 3 equations:
r + b -a = 10
a = 2r
b = 2r-15
----------------
Substitute for "a" and for "b" to get:
r + (2r-15) - (2r) = 10
r - 15 = 10
r = 25 cents (cost of one orange)
a = 2r = 50 cents (cost of one apple)
b = 2r-15 = 50-15 = 25 cents (cost of one banana)
=====================================================
Cheers,
Stan H.